OpenStudy (nevermind_justschool):

Variable y varies directly with variable x, and y = 4 when x = 7. What is the value of x when y = 12? I don't know, How can I get the answer?

OpenStudy (anonymous):

If i am interpretting the problem properly x=y

OpenStudy (anonymous):

okay like before y=kx since it is directly varying now put values of y and x to find k which is constant of proportionality 4=k7 4/7=k now y=4/7x 12=4/7x x=21 ans

OpenStudy (anonymous):

^

OpenStudy (nevermind_justschool):

Thank you so much, do you mind if I write this down for notes?

OpenStudy (anonymous):

of course use it in anyway you want.